# Universal framework for anisotropic nanoparticles with resonance laws and splitting

### v7

@@ Abstract
Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and mode response in subwavelength structures. Anisotropic nanoparticles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions. Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles. Closed-form solutions reveal axial-permittivity sum rules and material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns. Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss. Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory. This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, enabling a new generation of photonic devices with tunable multispectral response and directional emission.
Main
@@
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime, Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [23, 24, 25, 26, 27, 28, 29]. Analytic resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [30, 31, 32, 1, 33, 34, 35, 36, 37, 38, 39]. Remarkably, to date, the only structure in physics for which closed-form eigenmode response has been derived is the subwavelength isotropic sphere [40, 41, 32] (without residual one-dimensional integrals [30]). Recently, the emergence of anisotropic materials in optics [10, 42, 43, 11, 44] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [45, 46, 47, 48, 49, 50, 50, 51, 52, 53]. In particular, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and $\alpha$-molybdenum trioxide $(\alpha$-MoO3), making them the most promising among anisotropic particles for a broad range of nanotechnology applications [54, 11, 55, 17]. Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55]. However, whereas previous approaches have successfully treated isotropic particles, fundamental challenges arise when analyzing nanostructures with cartesian anisotropy—particularly in predicting higher-order mode behavior—which hinders their theoretical understanding. These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems. This motivates our exploration of a new theoretical framework. The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [11, 56] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [57, 58]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [11, 59, 60, 61, 55].
@@ Results
To allow several degrees of freedom of the permittivities in the different axes, we derive the eigenstates [41, 32] of uniaxial and biaxial particles in the quasistatic regime. Our starting point is solving Laplace’s equation, which reads for anisotropic media [67]
@@
where $\psi_{n}$ is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor $\overleftrightarrow{\epsilon_{n}}=\overleftrightarrow{\epsilon_{1n}}$ inside the inclusion and $\overleftrightarrow{\epsilon_{n}}=I\epsilon_{2}$ in the host medium ($I$ is the identity matrix). Note that the physical permittivity tensor $\overleftrightarrow{\epsilon_{1}}(\omega)$ generally depends on $\omega$ and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, $\boldsymbol{E}_{n}=\nabla\psi_{n}$ have to satisfy the boundary conditions at the particle interface. Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition. In Fig. 1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most prevalent types of particles.
@@
Conclusion
We derived the modes and eigenpermittivity relations of uniaxial and biaxial nanoparticles in the quasistatic regime as well as their radiation patterns. We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting. We also identified eigenfrequency degeneracy for uniaxial spheres, resulting in superimposed modes. Interestingly, as opposed to isotropic-material particles, high-order resonance frequencies in anisotropic particles are often lower than the dipole resonance. We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry. Finally, we derived the Q factors analytically and performed extensive full-wave simulations that highlighted the predictive power of the theory. Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [72, 73, 74, 75, 1]. The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers [59] with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59]. Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 76, 77]. Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
@@ Methods
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes $x,y$ and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of $11.5\mathrm{nm},$ and we situate an oscillating dipole at $\mathbf{r}_{0}=(22\,\mathrm{nm},0,22\,\mathrm{nm})$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$ We perform frequency domain simulations with a frequency sweep from $f=4.3652\cdot 10^{13}\,$(1/s) to $f=4.7626\cdot 10^{13}\,$(1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of $\hat{r}$ since we assume [67, 78, 32], and prove in the SM, that the mode excitation coefficient depends on $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$ so that we need to incorporate only $\partial\psi_{l}/\partial r,$ preserving the angular dependence in $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$. This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in Methods D. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by $\left\langle\psi_{k}|\psi_{0}\right\rangle=\int\theta_{1}\nabla\psi_{k}^{*}\cdot\nabla\psi_{0}d\mathbf{r^{\prime}},$ where $\theta_{1}$ is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the $x$ axis. In this case, we do not expect the $m=\pm(l-1)$ modes to show in the spectrum since in order to excite them a $z$ component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$ compared to the surface integral of $|\mathbf{E}|^{2}$ over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from $1/|\mathrm{resonance\,condition}(\omega)|^{2}=\mathrm{max}/2$ and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to $\epsilon_{x}=-1.5,$ is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by $\omega/\Delta\omega.$ We then plot for all the observed peaks $|\mathbf{E}|$ inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot $\int|\mathbf{E}|^{2}da$ for a dipole located at the same distance as the previous case and oriented along $x$ ($\mathbf{r}_{0}=(1,0,0)$31nm). As we predicted the $m=\pm(l-1)$ modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles. In Fig. 5 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode $\psi_{2}^{2}+\psi_{2}^{-2}$ with the resonance condition $\epsilon_{1x}=\epsilon_{1y}=-1.5$ outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with $R=46$ nm and $\epsilon_{1x}=-1.49,\epsilon_{1y}=-1.49,\epsilon_{1z}=1,$ excited by a current source similarly to Fig. 4 (top) in the radiation zone in the radiation zone $(r>2\lambda)$. Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere. As can be seen, the radiation-pattern results show very good agreement.
@@
One can express the eigenvalue in the denominator $s_{1xl}$ in terms of the physical parameter $\epsilon_{1z}$ using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs. 2 and 3). Note that the first sum includes only the $z$ dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper. It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.

### v8

# Universal framework for uniaxial and biaxial particles with resonance laws and splitting

@@ Abstract
Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and mode response in subwavelength structures. Uniaxial and biaxial particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions. Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of such nanoparticles. Closed-form solutions reveal axial-permittivity sum rules and material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns. Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss. Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles are consistent with the theoretical predictions and with our recently reported experimental observations. This framework unifies the understanding of anisotropic nanostructures across optics, magnetism, and thermal transport enabling a new generation of photonic devices with tunable multispectral response and directional emission.
Introduction
@@
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime, Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [23, 24, 25, 26, 27, 28, 29]. Analytic resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [30, 31, 32, 1, 33, 34, 35, 36, 37, 38, 39]. Remarkably, to date, the only structure in physics for which closed-form eigenmode response has been derived is the subwavelength isotropic sphere [40, 41, 32] (without residual one-dimensional integrals [30]). Recently, the emergence of anisotropic materials in optics [10, 42, 43, 11, 44] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [45, 46, 47, 48, 49, 50, 50, 51, 52, 53]. In particular, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and $\alpha$-molybdenum trioxide $(\alpha$-MoO3), making them the most promising among anisotropic particles for a broad range of nanotechnology applications [54, 11, 55, 17]. Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55]. However, whereas previous approaches have successfully treated isotropic particles, fundamental challenges arise when analyzing nanostructures with cartesian anisotropy—particularly in predicting higher-order mode behavior—hindering their theoretical understanding. These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, the distinct differential equations in each medium, and the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems. This motivates our exploration of a new theoretical framework. The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [11, 56] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [57, 58]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [11, 59, 60, 55].
Figure 1: Scope of applicability of the proposed framework
@@ Results
Closed-form eigenmodes and resonance conditions in uniaxial and biaxial spheres
To allow several degrees of freedom of the permittivities in the different axes, we utilize the definition of an eigenstate [41, 32] as a source-free electric potential in the quasistatic regime applied here to uniaxial and biaxial particles. Our starting point is solving Laplace’s equation, which reads for anisotropic media [66]
@@
where $\psi_{n}$ is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor $\overleftrightarrow{\epsilon_{n}}=\overleftrightarrow{\epsilon_{1n}}$ inside the inclusion and $\overleftrightarrow{\epsilon_{n}}=I\epsilon_{2}$ in the host medium ($I$ is the identity matrix). Note that the physical permittivity tensor $\overleftrightarrow{\epsilon_{1}}(\omega)$ generally depends on $\omega$ and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, $\boldsymbol{E}_{n}=\nabla\psi_{n}$ have to satisfy the boundary conditions at the particle interface. Once anisotropic Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition. In Fig. 2, left column, we present isotropic, uniaxial, and biaxial particles, which are the most prevalent types of particles.
@@
Analytic eigenmodes and resonance conditions in uniaxial and biaxial ellipsoids
@@
Discussion
We derived the modes and eigenpermittivity relations of uniaxial and biaxial nanoparticles in the quasistatic regime as well as their radiation patterns. We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting. We also identified eigenfrequency degeneracy for uniaxial spheres, resulting in superimposed modes. Interestingly, as opposed to isotropic-material particles, high-order resonance frequencies in anisotropic particles are often lower than the dipole resonance. We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry. Finally, we derived the Q factors analytically and performed extensive full-wave simulations that highlighted the predictive power of the theory. Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [71, 72, 73, 74, 1]. The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection. Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [59, 75, 76]. Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Finally, our results apply also to other fields of physics such as magnetism and heat conduction [56].
@@ Methods
We demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes $x,y$ and the permittivity model used in Figs. 3,4 [54]. The sphere has a radius of $11.5\mathrm{nm},$ and we situate an oscillating dipole at $\mathbf{r}_{0}=(22\,\mathrm{nm},0,22\,\mathrm{nm})$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$ We perform frequency domain simulations with a frequency sweep from $f=4.3652\cdot 10^{13}\,$(1/s) to $f=4.7626\cdot 10^{13}\,$(1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of $\hat{r}$ since we assume [66, 77, 32], and prove in the SM, that the mode excitation coefficient depends on $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$ so that we need to incorporate only $\partial\psi_{l}/\partial r,$ preserving the angular dependence in $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$. This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in Methods D. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by $\left\langle\psi_{k}|\psi_{0}\right\rangle=\int\theta_{1}\nabla\psi_{k}^{*}\cdot\nabla\psi_{0}d\mathbf{r^{\prime}},$ where $\theta_{1}$ is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the $x$ axis. In this case, we do not expect the $m=\pm(l-1)$ modes to show in the spectrum since in order to excite them a $z$ component of the dipole location is required. In Fig. 5 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$ compared to the surface integral of $|\mathbf{E}|^{2}$ over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from $1/|\mathrm{resonance\,condition}(\omega)|^{2}=\mathrm{max}/2$ and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to $\epsilon_{x}=-1.5,$ is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by $\omega/\Delta\omega.$ We then plot for all the observed peaks $|\mathbf{E}|$ inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot $\int|\mathbf{E}|^{2}da$ for a dipole located at the same distance as the previous case and oriented along $x$ ($\mathbf{r}_{0}=(1,0,0)$31nm). As we predicted the $m=\pm(l-1)$ modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles. In Fig. 6 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode $\psi_{2}^{2}+\psi_{2}^{-2}$ with the resonance condition $\epsilon_{1x}=\epsilon_{1y}=-1.5$ outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with $R=46$ nm and $\epsilon_{1x}=-1.49,\epsilon_{1y}=-1.49,\epsilon_{1z}=1,$ excited by a current source similarly to Fig. 5 (top) in the radiation zone in the radiation zone $(r>2\lambda)$. Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere. As can be seen, the radiation-pattern results show very good agreement.
@@
One can express the eigenvalue in the denominator $s_{1xl}$ in terms of the physical parameter $\epsilon_{1z}$ using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs. 3 and 4). Note that the first sum includes only the $z$ dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper. It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.